

********************************************************************
UARS WEEKLY STATUS REPORT - 22 JUL 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was turned on July 18. Since then, it has successfully gathered
daily mid (1.1 nm) resolution, MgII, and continuous monitoring scans
for time periods with UARS telemetry or tape recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 July 2005 (Orbit 75678, GMT Day 190) through
15 July 2005 (Orbit 75783, GMT Day 196).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 194/1658 75748 POWER ON
HALOE 194/1801 75749 POWER OFF
SOLSTICE 195/1258 75761 DAY ONLY TO FULLTIME
SOLSTICE 196/1436 75777 POWER ON TO FULLTIME
A Forward-to-Reverse Yaw Around maneuver (number 141) was successfully
executed in Orbit 75702 beginning 191/ 14:54:22Z and completing at
191/ 15:31:32Z on 10 July 2005. All instruments were in their off or
standby mode (except for SOLSTICE in day-only). Due to the SA
position at 265 degrees, the effective night length of the spacecraft
after the maneuver was approximately 59.80 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 25.12 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 25.40 and 14.45 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 9
August 2005 (DOY 222).
The beta angle ranged between the angles of 05.0 degrees to 19.0
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 37 degrees on 24 July 2005 (DOY 205).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 4.84 to 4.48 degrees C. Additional
battery performance data for each battery on 08 July 2005 and 15 July
2005 are provided in Appendix D.
The clock error ranged from 10.1 to -10.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
HALOE = 02
SOLSTICE = 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
On Monday 11 July 2005 the results of the senior review at NASA HQ
were received. The direction given is that UARS will fly with HALOE
observations only until the end of the calendar year, followed by the
decommissioning of the UARS spacecraft in the new year. Final
calibrations for SUSIM will be performed before they are turned off
completely. UARS will fly with HALOE only after the clock rollover.
The UARS OBC clock rollover effort continued this week with more
discussions regarding contingency plans as well as an analysis of the
HGA slew angles during the uplink. The FOT received the preliminary
patch for the Eclipse ground system 11 July 2005 and testing has been
conducted. Several mnemonics still showed the wrong time. Raytheon
has been made aware of these findings and the FOT is awaiting another
iteration of the patch.
MEETINGS
The FOT attended the UARS MSR on 12 July 2005 to discuss the current
status of the spacecraft and future plans for the rest of the UARS
mission.
The FOT also attended the weekly clock rollover meeting to review
activities that need to be finished by 02 August 2005. So far the
project is still on schedule a few items still need to be completed
before the FOT is go for uplink.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 9 August 2005 (DOY 222).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 15 July 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
75678 190/0024 05-211 MPS VT L5 TO VT L6
75701 191/1357 05-212 MACS YAW FWD TO REV
75725 193/0343 05-213 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
75746 194/1301 05-214 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
75748 194/1658 05-215 HALOE HALOE POWER ON PT 1
75749 194/1801 05-216 HALOE HALOE POWER OFF
75761 195/1258 05-217 SOLSTICE SOLSTICE FROM DAY-ONLY TO FULLTIME
75767 195/2238 05-218 SOLSTICE ATTEMPTED INSTR RE-BOOT
75777 196/1436 05-219 SOLSTICE POWER ON - FULLTIME OPS
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 15 July 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 15 July 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
75725 193/0408 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
75746 194/1301 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 15 July 2005
08 July 2005 - GMT Day 189 Beta = 09.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 22.4/ 0.0 -3.68/-4.94 +00.0/+00.0 05.8 0.00 0.0
2 0.0/-112.0 +6.44/+5.15 +31.2/-17.6 25.3 1.02 21.5
3 212.8/-700.0 +2.29/+1.34 +20.8/-09.6 25.1 1.04 13.4
15 July 2005 - GMT Day 196 Beta = 19.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 39.2/ 11.2 -3.05/-4.31 +00.0/+00.0 05.8 0.00 0.0
2 0.0/-134.4 +8.07/+7.08 +32.0/-18.0 26.4 1.04 22.5
3 112.0/-700.0 +3.23/+2.60 +18.8/-10.0 26.2 1.06 12.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-NONE THIS REPORT PERIOD
APPENDIX E
UARS ANOMALY REPORTS
Week ending 15 July 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-034 FOT 75686 190/1258 40062 White Sands UPLINK Problem
05-032 FOT 75693 191/0057 30832 Generic Late Acquisition
05-035 SOLSTICE 75767 195/2238 N/A SOLSTICE in incorrect
state
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-033 FOT 75656 188/1303 40038 Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-032 FOT See List July 05 30832 Generic Late Acquisition
05-034 FOT 75686 190/1258 40062 White Sands UPLINK Problem
05-035 SOLSTICE 75767 195/2238 N/A SOLSTICE in incorrect
state